For this, use compounded interest rate unless it explicitly requires using simple interest.

You would like an initial investment of $9,000 invested now to grow to $20,000 in 10 years. What interest rate (to the nearest tenth of a percent) must you earn to achieve your goal?
Answer #1

ANSWER:

Future value = 20,000

present value = 9,000

n =10 years

i = ?

fv = pv * (1 + i) ^ n

20,000 = 9,000 * (1 + i) ^ 10

20,000 / 9,000 = (1 + i) ^ 10

2.22 = (1 + i) ^ 10

(2.22) ^ 1/10 = 1 + i

1.083017 = 1 + i

i = 1.083017 - 1

i = 0.083125 or 8.31%

so the interest rate to achieve the goal is 8.31%